OpenStudy (anonymous):

Can someone help me please? A wooden pyramid, 12 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 5 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (anonymous):

Do you know what the volume of a pyramid is?

OpenStudy (anonymous):

1/3 (b)(h)?

OpenStudy (anonymous):

Yep. So in the first case, what would the volume be? You to to find the ration you divide the new value by the old value right?

OpenStudy (anonymous):

To find the ratio*

OpenStudy (anonymous):

Well I know the height, but what's the base?

OpenStudy (anonymous):

No need. You got \[v=\frac{ 1 }{ 3 }A*h\]

OpenStudy (anonymous):

h=12.

OpenStudy (anonymous):

Therefore: \[v=\frac{ 1 }{3 } * A *12\] = \[V=4A\]

OpenStudy (anonymous):

Now we are given the ara of the base was increased by a factor of 5. In other words, it was multiplied by 5.

OpenStudy (anonymous):

So, 20?

OpenStudy (anonymous):

Ratios mean you divide, not multiply.

OpenStudy (anonymous):

Now I'm confused :(

OpenStudy (anonymous):

About?

OpenStudy (anonymous):

I thought you said it was multiplied by 5?

OpenStudy (anonymous):

Haha, yeah for the new volume. But when you find a ration you divide the new value by the old value.

OpenStudy (anonymous):

ratio*

OpenStudy (anonymous):

So... what is the ratio of the volumes of the 2 pyramids? o.o

OpenStudy (anonymous):

Lol, I don't want to give you an answer :P . I would rather you understand it.

OpenStudy (anonymous):

Allright better explanation.

OpenStudy (anonymous):

It's a 1/5 ratio right?

OpenStudy (anonymous):

Because we scale the pyramid by a factor of 5.

OpenStudy (anonymous):

1:5 ratio sorry.

OpenStudy (anonymous):

Ignore all the other parts.

OpenStudy (anonymous):

I'm seriously so confused it's not even funny! Lol okay so the dimensions are increased by a facor of 5. So you're saying the ratio of the 2 pyramids is 1:5?

OpenStudy (anonymous):

Ahh but remember, since this is volume, we cube that 5.

OpenStudy (anonymous):

1/5^3

OpenStudy (anonymous):

or 1:125 which can also be 1/125.

OpenStudy (anonymous):

My head is starting to hurt..so I cube the ratio 1:5? WHAT? Lol you have to break it down for me all at once. I don't do step by step :b

OpenStudy (anonymous):

Allright.

OpenStudy (anonymous):

The scale factor of the Pyramid is 1:5 right? This can be re-written as 1/5 . So the ratio would be 1:5 but since we are dealing with volumes, We have to cube that ratio or in other words, 1/5^3 or 1/125. We can say the New pyramid is 125x more greater than the old pyramid.

OpenStudy (anonymous):

Is that the answer or are you waiting for me to answer? Lol I don't know how to do this.

OpenStudy (anonymous):

What don't you understand?

OpenStudy (anonymous):

I want to try and help :c .

OpenStudy (anonymous):

I'm confused on the entire thing lol, I mean I'd like to understand how to get the right answer so I can fully explain my answer in complete sentences but at this point my head is starting to hurt and I just want to finish it. Lol

OpenStudy (dumbcow):

initial dimensions; height = 12 let the square base have dimensions of 1 so initial volume = 4 new dimensions, multiply lengths by 5 height = 60 square base have sides of 5 .... so area of 25 new volume = (1/3)(25)(60) = 500 volume ratio = 500/4 = 125 does that help?

OpenStudy (dumbcow):

in general, the rule is that the volume ratio = (length ratio)^3 --> 5^3 = 125

OpenStudy (anonymous):

Yes! Thank you both for helping me.. sorry I get confused so easily! I appreciate you both! :)

OpenStudy (dumbcow):

:)